1600 | 1700 | 1800 | 1900 | 2000 | 2100 | 2200 | 2300 | Total Time | |
---|---|---|---|---|---|---|---|---|---|
Acceptable Beam in Use | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.54 | 0.79 | 0.20 | 1.53 |
Beam Available but Not in Use | 0.89 | 0.97 | 0.91 | 0.93 | 0.25 | 0.02 | 0.10 | 0.80 | 4.87 |
Beam Not Available or Unacceptable | 0.11 | 0.03 | 0.09 | 0.07 | 0.75 | 0.44 | 0.12 | 0.00 | 1.60 |
Accelerator Con- figuration Change | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Experiment Ready | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.75 | 1.00 | 0.25 | 2.00 |
Planned Config- uration Change | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Unplanned Experiment Down | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 0.25 | 0.00 | 0.75 | 6.00 |
Hall not ready reasons | |||||||||
Target | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Data Acq | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Other | 0.75 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.75 |
HRSL | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
HRSR | 0.25 | 1.00 | 1.00 | 1.00 | 1.00 | 0.25 | 0.00 | 0.75 | 5.25 |
HALL_OFF | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Comments: 16:30 access to reset right arm magnet controls 16:45 rebooted the magnet control ioc of the right arm 16:46 right arm septum tripped 17:00 Brian fixed the sync problem in the right arm 17:30 both dipoles have tripped 19:45 no hope to fix right dipole today - take data with one arm. end of access. 20:45 injector problem no beam yet 21:15 beam back 5 muA 21:20 beam at 20 muA 21:25 beam at 40 muA 21:35 requested 60 muA 21:50 MCC called because they cannot do better than 40 muA because of trip in the ion chamber in the left dump 22:10 after David's intervention they raise the trip limit and got us to 60 muA 22:30 adjust detector HV to accouont for these conditions. 5655 - production run with 1 arm 60 muA. 23:00 we will turn the right septum on. ask for beam off and wait that the temperature of the coil goes below the limit of 7.53K. 23:10 start ramping up the current in the R septum to 316.92 and set the L septum at 323.61 23:45 beam back 5656 - production run with 1 arm 58 muA
Data Entry by: E. Burtin